# Efficient Dynamic Scene Editing via 4D Gaussian-based Static-Dynamic Separation

### v1

@@ 4 Method
In this section, we explain our efficient dynamic scene editing method illustrated in Fig. 3. (Sec. 4.1) We first train dynamic scenes for editing targets by optimizing the 4D Gaussian Splatting (4DGS)  [60]. (Sec. 4.2) Motivated by the static-dynamic separability of Hexplane [15, 5]-based 4DGS, we initially focus on editing the static canonical 3D Gaussians  [23] to efficiently edit the dynamic scene. (Sec. 4.3) To mitigate overfitting issues that may arise during the 3D Gaussian editing process and refine the motion artifacts in the dynamic scene, we introduce a temporal refinement stage using score distillation.
@@ 4.1 Optimizing 4D Gaussians for Target Scenes
To perform editing, we first require the 4D Gaussian representation of the target dynamic scene. To obtain this, we optimize the 4D Gaussians and use it for editing. Specifically, we use a dynamic scene dataset [27] composed of multi-camera captured videos, which can be represented as a set of images $\{I_{M,t}\}$ in which $M$ denotes the camera matrix and $t$ denotes the timestep within the videos. We synthesize images $\hat{I}_{M,t}$ by rendering the randomly initialized dynamic scene $\left\{\mathcal{G}_{\text{canon}}^{\text{init}},\mathcal{E}^{\text{init}}(\mathcal{G}_{\text{canon}},t),\mathcal{D}^{\text{init}}\right\}$. Then we calculate the RGB L1 loss against the corresponding dataset image $I_{M,t}$, training the dynamic scene through this process. We also applied a grid-based total variational loss [15, 5, 53, 14], $\mathcal{L}_{\text{TV}}$ to ensure smoothness in the deformation field output along the timesteps. Note that, as our editing method is highly dependent on the quality of the target dynamic scene, incorporating such regularization loss is helpful. The entire loss function used for 4DGS training is: $\mathcal{L}_{\text{4DGS}}=|\hat{I}_{M,t}-I_{M,t}|+\mathcal{L}_{\text{TV}}$.
As a result, we obtain the optimized 4D Gaussians $\left\{\mathcal{G}_{\text{canon}}^{\text{opt}},\mathcal{E}^{\text{opt}}(\mathcal{G}_{\text{canon}},t),\mathcal{D}^{\text{opt}}\right\}$, which represent the editing target dynamic scene. The static component $\mathcal{G}_{\text{canon}}^{\text{opt}}$ serves as the main editing target in Sec. 4.2 to Sec. 4.3. To resolve the misalignment between the edited static component $\mathcal{G}_{\text{canon}}^{\text{edit}}$ and the dynamic components $\left\{\mathcal{E}^{\text{opt}}(\mathcal{G}_{\text{canon}},t),\mathcal{D}^{\text{opt}}\right\}$, we apply score [41]-based temporal refinement in Sec. 4.3. A more detailed training setup follows [60].
@@ 4.2 Stage 1: Efficient Dynamic Scene Editing with Static 3D Gaussians
In Sec. 4.1, we obtained the optimized canonical 3D Gaussians $\mathcal{G}_{\text{canon}}^{\text{opt}}$, which models the explicit appearance and geometry of a 3D scene, along with the optimized dynamic components $\mathcal{E}^{\text{opt}}(\mathcal{G}_{\text{canon}},t)$ and $\mathcal{D}^{\text{opt}}$. For efficient dynamic scene editing, we perform editing only on $\mathcal{G}_{\text{canon}}^{\text{opt}}$, which is minimal but sufficient static information for visual editing of the dynamic scene as shown in Fig. 3.
To obtain supervision images for editing the $\mathcal{G}_{\text{canon}}^{\text{opt}}$, we extract a subset of multiview images fixed at the initial timestep and edited them using InstructPix2Pix [4]. Subsequently, we edit optimized canonical 3D Gaussians $\mathcal{G}_{\text{canon}}^{\text{opt}}$ with an L1 RGB loss supervising the edited images. Compared to the latest 4D editing method Instruct 4D-to-4D [37], which requires editing (number of video timesteps $T$ number of cameras $\mathcal{M}$) images through iterative dataset updates, our approach significantly reduces the computation required to address the editing of the dynamic scene. Moreover, this approach enables fast shifts to the edited result, regardless of the number of timesteps $T$ of the dynamic scene. After completing the 3D Gaussian editing process, we can obtain a pseudo-edited dynamic scene $\left\{\mathcal{G}_{\text{canon}}^{\text{edit}},\mathcal{E}^{\text{opt}}(\mathcal{G}_{\text{canon}},t),\mathcal{D}^{\text{opt}}\right\}$, which is obtained by simply recombining the edited canonical 3D Gaussians $\mathcal{G}_{\text{canon}}^{\text{edit}}$ with the original Gaussian deformation field $\left\{\mathcal{E}^{\text{opt}}(\mathcal{G}_{\text{canon}},t),\mathcal{D}^{\text{opt}}\right\}$.
@@ 4.3 Stage 2: Refinement using Score Distillation for Temporal Alignment
After the first editing stage proposed in Sec. 4.2, the pseudo-edited dynamic scene $\left\{\mathcal{G}_{\text{canon}}^{\text{edit}},\mathcal{E}^{\text{opt}}(\mathcal{G}_{\text{canon}},t),\mathcal{D}^{\text{opt}}\right\}$ exhibits severe motion artifacts, as shown in Fig. 4 (a). The main reason is that during the 3D Gaussian editing process, the positions $p$ of Gaussian primitives in $\mathcal{G}_{\text{canon}}^{\text{opt}}$ shift slightly, causing the queried embedding voxel features $f_{h}$ to differ from the original dynamic scene. Moreover, only the Spherical Harmonics (SH) colors on the surface visible at the initial timestep are updated. As a result, if the Gaussian primitives in pseudo-edited $\mathcal{G}_{\text{canon}}^{\text{edit}}$ rotate at later timesteps, unedited SH values that were previously hidden may become exposed, leading to artifacts. Therefore, we introduce a temporal refinement stage to resolve the misalignment between the original deformation field $\left\{\mathcal{E}^{\text{opt}}(\mathcal{G}_{\text{canon}},t),\mathcal{D}^{\text{opt}}\right\}$ and the edited canonical 3D Gaussians $\mathcal{G}_{\text{canon}}^{\text{edit}}$.
To perform the refinement stage efficiently without editing additional dataset images, we employed the score distillation mechanism  [41]. Since the dynamic scene is also synthesized from multiple 2D images, the prior of the 2D diffusion model can be distilled into the 4D dynamic scene representation. The editing process can be continued using the noise prediction loss (i.e. score) obtained from each IP2P inference as Eqs. 1 and 2. Since we just use score distillation for editing refinement rather than generation or editing from scratch, this stage can be completed with a smaller number of iterations. Consequently, our approach is relatively less affected by inherent issues of Score Distillation Sampling (SDS), such as the Janus problem [50].
Similar to Sec. 4.2, we applied Coherent-IP2P with the diffusion prior $\theta$ and observed that it reduces blurring effects and enhances qualitative performance compared to the original IP2P. At each refinement iteration, we rendered $B$ images of the pseudo-edited dynamic scene $\tilde{I}=\left\{\hat{I}_{i}=S(M_{i},\mathcal{G}_{\text{def},t_{i}}^{\text{edit}}+\Delta G_{t,i})\right\}_{i=1}^{B}$ using random camera matrices $\left\{M_{i}\right\}_{i=1}^{B}$ and random timesteps $\left\{t_{i}\right\}_{i=1}^{B}$ as input for Coherent-IP2P, where $S$ denotes the rendering process of the 3DGS (subscripts $M$ and $t$ on $\hat{I}$ omitted for simplicity). The SDS loss we used to obtain refined 3D Gaussians $\mathcal{G}_{\text{canon}}^{\text{ref}}$ is as follows:
@@
, where $\tilde{t}$ is diffusion timestep, $\epsilon$ is diffusion noise, $c_{I}=\left\{I_{i}\right\}_{i=1}^{B}$ is original dataset images, $c_{T}$ is user instruction, $s_{I}$ and $s_{T}$ are Classifier-Free-Guidance [19] scale for $c_{I}$ and $c_{T}$, $\epsilon_{\theta}(\tilde{I},c_{I},c_{T})$ is Coherent-IP2P denoiser networks including VAE [25]. This score-based guidance encourages a set of rendered 2D images from the pseudo-edited 4D Gaussians at arbitrary timesteps $\tilde{I}$ to resemble the edited images that IP2P would generate based on the $c_{I}$ and $c_{T}$, thereby effectively refining motion artifacts. As a result, we obtain refined canonical 3D Gaussians $\mathcal{G}_{\text{canon}}^{\text{ref}}$ that aligns well with the original deformation field $\left\{\mathcal{E}^{\text{opt}}(\mathcal{G}_{\text{canon}},t),\mathcal{D}^{\text{opt}}\right\}$ while maintaining the edited appearance. After the refinement stage, we finally obtain a completely edited 4D dynamic scene represented by 4D Gaussians as $\left\{\mathcal{G}_{\text{canon}}^{\text{ref}},\mathcal{E}^{\text{opt}}(\mathcal{G}_{\text{canon}},t),\mathcal{D}^{\text{opt}}\right\}$.
Figure 4: Effectiveness of score-based temporal refinement: Proposed score-based temporal refinement effectively resolves the misalignment between the canonical 3D Gaussians and the original deformation field that arises during the 3D Gaussian editing process. Without requiring additional 2D image updates, this process efficiently completes the dynamic scene editing within a few hundred iterations.
@@ 5.1 Experimental Setup
We utilize the real-world multiview video dataset, DyNeRF [27] to train and edit the 4D dynamic scene. The DyNeRF dataset includes six 10-second video sequences captured at 30 fps by 15 to 20 cameras with face-forward perspective. Therefore, extracting each timestep of the video as an image results in thousands of images.
@@
We conduct a qualitative and quantitative comparison with Instruct 4D-to-4D [37], the only prior work addressing instruction-guided 4D dynamic scene editing. Instruct 4D-to-4D utilizes NeRFPlayer [52] as its backbone 4D representation and employs an iterative dataset update method, editing all 2D images used in synthesizing the dynamic scene. It utilizes optical flow-based warping [55] and depth-based warping to ensure consistency across all edited 2D images. To alleviate the time-consuming dataset update process, Instruct 4D-to-4D uses two GPUs in parallel: one for the dataset updating (i.e. editing) thread and the other for the dynamic scene editing thread.
@@
In Sec. 4.1, we follow the experimental settings of [60]. Throughout the experiments utilizing InstructPix2Pix [4], we set the CFG [19] scales for image condition and text instruction to 1.2 and 10.5, respectively. In the 3D Gaussian editing stage (Sec. 4.2), 800 to 1000 iterations were performed, depending on the editing style. For the score-based [41] refinement stage (Sec. 4.3), an average of 500 iterations is sufficient to successfully complete the dynamic scene editing. All experiments were conducted using a single NVIDIA A40 GPU.
@@ 5.2 Results
We measure PSNR, SSIM [58], and LPIPS [66] between the rendered images of the edited dynamic scene and the edited 2D multiview images used in the 3D Gaussian editing stage. Note that, as the temporal refinement stage can lead to an unfair evaluation of these metrics since the score loss does not provide direct supervision for the edited 2D images, we measured them by rendering the pseudo-edited dynamic scenes instead. To evaluate the refined dynamic scene, we measure CLIP [43] similarity to evaluate how well the input instruction was followed. Note that, the refined dynamic scenes exhibit a higher CLIP similarity compared to the pseudo-edited dynamic scenes.
A quantitative comparison between our method and the baseline is presented in Tab. 1. Our edited dynamic scene outperforms the baseline across all metrics, demonstrating that our method can produce higher-quality edited dynamic scenes while better following the user instruction. Tab. 2 presents the results comparing the efficiency of the baseline and the proposed method. In terms of editing time, our method completes editing 2–3 times faster than the baseline, despite using only a single GPU compared to the baseline’s use of two identical GPUs. This difference in editing time is expected to increase as the number of timesteps covered by the dynamic scene grows. These results demonstrate that the proposed method, which ensures efficiency by focusing on the static-dynamic separability of 4DGS and employing score-based temporal refinement, performs dynamic scene editing much more efficiently without the need for extensive time-consuming dataset updates.
@@
Ours 27.22 0.856 0.250 0.256
Table 1: Quantitative comparison of editing quality: Comparison of performance metrics between Instruct 4D-to-4D (denoted as I4D24D) and our method for different instructions. Higher values are better for PSNR, SSIM, and CLIP similarity, while lower values are better for $\text{LPIPS}_{\text{VGG}}$.
@@
Ours 1 GPU 40 minutes
Table 2: Quantitative comparison of editing efficiency: Compared to the baseline, our method can complete dynamic scene editing in a very short time using only a single GPU, with this advantage becoming more pronounced as the number of timesteps in the dynamic scene increases.
Figure 5: Qualitative results across various editing styles: We present the editing results for the cook_spinach, flame_steak, and coffee_martini scenes from DyNeRF. Our method successfully generates high-quality edited dynamic scenes that closely follow the given various user instructions.
@@
The following ablation study explores the importance of each design choice in the proposed method, highlighting their contributions to efficiency and quality. The qualitative and user study results are in Fig. 8. We conducted a user study to evaluate the results of our ablation study. We recruited 50 random participants of varying genders and ages and collected a total of 50 preference-ranking responses on the editing quality of videos generated by the four methods mentioned in Fig. 8. As shown in the pie charts in Fig. 8, our method demonstrates the highest preference.
@@
We first compared the results of dynamic scene editing performed solely with score-based editing, without engaging in 3D Gaussian editing (denoted as “Fully SDS”). As shown in Fig. 8, editing with only SDS loss preserved smooth motion but failed to achieve sufficient alignment with the instruction, resulting in low-fidelity output. In contrast, by using the 3D Gaussian editing stage, our method secured high fidelity while achieving effective motion refinement in the temporal refinement process without compromising fidelity or quality. This demonstrates that providing direct supervision with the edited 2D images through the 3D Gaussian editing stage is essential for ensuring editing fidelity and quality.
@@
To mitigate inherent issues of SDS such as the Janus problem [50] and to obtain stable editing guidance for the score-based temporal refinement, we utilized Coherent-IP2P. To verify the effectiveness of this choice, we conducted an experiment comparing the results by refining the pseudo-edited dynamic scene with the original IP2P (denoted as “Refine w/ original IP2P”). Using the original IP2P for temporal refinement results in severe visual artifacts and low-quality outputs, whereas Coherent-IP2P preserves details and retains the semantics of the original scene. This demonstrates that Coherent-IP2P mitigates noisy guidance and blurry artifacts by enabling information sharing among images within the same batch.
Our final method (denoted as “Ours (w/o refine $\{\mathcal{E},\mathcal{D}\}$)”) refines only the static 3D Gaussians in the refinement stage, excluding the deformation field. This approach was chosen because refining the deformation field with Coherent-IP2P, which is not a video diffusion model, results in temporal inconsistencies and motion artifacts, as observed in “Ours (w/ refine $\{\mathcal{E},\mathcal{D}\}$)”. (See the supplementary video file for a closer view of the qualitative results.)
Figure 6: Qualitative comparison of visual quality: As shown in the zoomed-in images, the baseline exhibits blurred areas while our method produces clear results, accurately capturing fine details such as clothing folds and object contours. This demonstrates both the performance of our dynamic scene representation 4DGS and the stability of our dynamic scene editing method.
@@
Figure 7: Qualitative comparison of temporal consistency: As shown in the highlighted, zoomed-in areas, the baseline exhibits a flickering effect as the timestep progresses. In contrast, our method avoids such artifacts by editing only the static information and applying score-based temporal refinement.
@@
Figure 8: Ablation study of the dynamic scene editing method: Each segment of the pie chart represents the proportion of responses indicating the preference rank for the above result. Only the proposed method (denoted as “Ours (w/o refine $\{\mathcal{E},\mathcal{D}\}$)”) successfully generates an appropriately edited dynamic scene, achieving the highest preference score in the user study. This demonstrates the effectiveness and validity of our dynamic scene editing method.
6 Conclusion

### v2

@@ 4 Method
In this section, we present Instruct-4DGS, our proposed method for efficient dynamic scene editing, as illustrated in Fig. 3. (Sec. 4.1) We first train dynamic scenes for editing targets by optimizing the 4D Gaussian Splatting (4DGS)  [63]. (Sec. 4.2) Motivated by the static-dynamic separability of Hexplane [15, 5]-based 4DGS, we initially focus on editing the static canonical 3D Gaussians  [24] to efficiently edit the dynamic scene. (Sec. 4.3) To mitigate overfitting issues that may arise during the 3D Gaussian editing process and refine the motion artifacts in the dynamic scene, we introduce a temporal refinement stage using score distillation.
@@ 4.1 Optimizing 4D Gaussians for Target Scenes
Our dynamic scene editing requires a 4D Gaussian representation of the target dynamic scene. To obtain this, we optimize the 4D Gaussians and use it for editing. Specifically, we use dynamic scene datasets [28, 50] composed of multi-camera captured videos, which can be represented as a set of images $\{I_{M,t}\}$ in which $M$ denotes the camera matrix and $t$ denotes the timestep within the videos. We synthesize images $\hat{I}_{M,t}$ by rendering the randomly initialized dynamic scene $\left\{\mathcal{G}_{\text{canon}}^{\text{init}},\mathcal{E}^{\text{init}}(\mathcal{G}_{\text{canon}},t),\mathcal{D}^{\text{init}}\right\}$. Then we calculate the RGB L1 loss against the corresponding dataset image $I_{M,t}$, training the dynamic scene through this process. We also apply a grid-based total variational loss [15, 5, 56, 14], $\mathcal{L}_{\text{TV}}$ to enforce smoothness in the deformation field output along the timesteps. Note that, as our editing method is highly dependent on the quality of the target dynamic scene, incorporating such regularization loss is helpful. The entire loss function used for 4DGS training is: $\mathcal{L}_{\text{4DGS}}=|\hat{I}_{M,t}-I_{M,t}|+\mathcal{L}_{\text{TV}}$.
As a result, we obtain the optimized 4D Gaussians $\left\{\mathcal{G}_{\text{canon}}^{\text{opt}},\mathcal{E}^{\text{opt}}(\mathcal{G}_{\text{canon}},t),\mathcal{D}^{\text{opt}}\right\}$, which represent the editing target scene. The static component $\mathcal{G}_{\text{canon}}^{\text{opt}}$ serves as the main editing target in Sec. 4.24.3. In Sec. 4.2, we edit the static component by modifying only the images corresponding to the first timestep, ensuring efficient editing. In Sec. 4.3, we refine the edited static component $\mathcal{G}_{\text{canon}}^{\text{edit}}$ to better align with the original deformation field $\left\{\mathcal{E}^{\text{opt}}(\mathcal{G}_{\text{canon}},t),\mathcal{D}^{\text{opt}}\right\}$ using score-based temporal refinement, mitigating potential motion artifacts. A more detailed training setup follows [63].
@@ 4.2 Stage 1: Efficient Dynamic Scene Editing with Static 3D Gaussians
In Sec. 4.1, we obtained the optimized canonical 3D Gaussians $\mathcal{G}_{\text{canon}}^{\text{opt}}$, which models the explicit appearance and geometry of a 3D scene, along with the optimized dynamic components $\mathcal{E}^{\text{opt}}(\mathcal{G}_{\text{canon}},t)$ and $\mathcal{D}^{\text{opt}}$. For efficient dynamic scene editing, we perform editing only on $\mathcal{G}_{\text{canon}}^{\text{opt}}$, which is minimal but sufficient information for visual editing of the dynamic scene as shown in Fig. 3.
To generate supervision images for editing the $\mathcal{G}_{\text{canon}}^{\text{opt}}$, we extract a subset of multiview images fixed at the initial timestep and then edit them using InstructPix2Pix [4]. Subsequently, we edit optimized canonical 3D Gaussians $\mathcal{G}_{\text{canon}}^{\text{opt}}$ with an L1 RGB loss supervising the edited images. Compared to the latest 4D editing method Instruct 4D-to-4D [38], which requires editing $T\!\times\!\mathcal{M}$ images—where $T$ is the number of video timesteps and $\mathcal{M}$ is the number of cameras—through iterative dataset updates, our approach significantly reduces the computation required to address the editing of the dynamic scene. Moreover, this approach allows rapid transitions to the edited result, regardless of the number of timesteps $T$ of the dynamic scene. After completing the 3D Gaussian editing process, we obtain a pseudo-edited dynamic scene $\left\{\mathcal{G}_{\text{canon}}^{\text{edit}},\mathcal{E}^{\text{opt}}(\mathcal{G}_{\text{canon}},t),\mathcal{D}^{\text{opt}}\right\}$, which is obtained by simply recombining the edited canonical 3D Gaussians $\mathcal{G}_{\text{canon}}^{\text{edit}}$ with the original Gaussian deformation field $\left\{\mathcal{E}^{\text{opt}}(\mathcal{G}_{\text{canon}},t),\mathcal{D}^{\text{opt}}\right\}$.
@@ 4.3 Stage 2: Refinement using Score Distillation for Temporal Alignment
After the first editing stage proposed in Sec. 4.2, the pseudo-edited dynamic scene $\left\{\mathcal{G}_{\text{canon}}^{\text{edit}},\mathcal{E}^{\text{opt}}(\mathcal{G}_{\text{canon}},t),\mathcal{D}^{\text{opt}}\right\}$ exhibits severe motion artifacts, as shown in Fig. 4 (a). The primary cause is the slight shift in the positions $p$ of Gaussian primitives in $\mathcal{G}_{\text{canon}}^{\text{opt}}$ during the 3D Gaussian editing process, which results in discrepancies between the queried embedding voxel feature $f_{h}$ and those of the original dynamic scene. Moreover, only the Spherical Harmonics (SH) colors on the surface visible at the initial timestep are updated. As a result, if the Gaussian primitives in pseudo-edited $\mathcal{G}_{\text{canon}}^{\text{edit}}$ rotate at later timesteps, unedited SH values that were previously hidden may become exposed, leading to artifacts. Therefore, we introduce a temporal refinement stage to resolve the misalignment between the original deformation field $\left\{\mathcal{E}^{\text{opt}}(\mathcal{G}_{\text{canon}},t),\mathcal{D}^{\text{opt}}\right\}$ and the edited canonical 3D Gaussians $\mathcal{G}_{\text{canon}}^{\text{edit}}$.
To perform the refinement stage efficiently without editing additional dataset images, we employ the score distillation mechanism  [43]. Since the dynamic scene is edited using multiple 2D images generated by IP2P, the prior of the 2D diffusion model (i.e., IP2P) can be distilled into the 4D dynamic scene. The editing process can be continued using the noise prediction loss (i.e. score) obtained from each IP2P inference as Eqs. 1 and 2. Since we just use score distillation for editing refinement rather than generation or editing from scratch, this stage can be completed with a smaller number of iterations. Consequently, our approach is relatively less affected by inherent issues of Score Distillation Sampling (SDS), such as Janus problem [53].
Similar to Sec. 4.2, we apply Coherent-IP2P with the diffusion prior $\theta$ and observe that it reduces blurring effects and enhances qualitative performance compared to the original IP2P. At each refinement iteration, we rendered $B$ images of the pseudo-edited dynamic scene $\tilde{I}=\left\{\hat{I}_{i}=S(M_{i},\mathcal{G}_{\text{def},t_{i}}^{\text{edit}})\right\}_{i=1}^{B}$ using random camera matrices $\left\{M_{i}\right\}_{i=1}^{B}$ and random timesteps $\left\{t_{i}\right\}_{i=1}^{B}$ as input for Coherent-IP2P, where $S$ denotes the rendering process of the 3DGS (subscripts $M$ and $t$ on $\hat{I}$ omitted for simplicity). We optimize $\mathcal{G}_{\text{canon}}^{\text{edit}}$ using the following SDS loss to obtain the refined 3D Gaussians $\mathcal{G}_{\text{canon}}^{\text{ref}}$:
@@
, where $\tilde{t}$ is diffusion timestep, $\epsilon$ is diffusion noise, $c_{I}=\left\{I_{i}\right\}_{i=1}^{B}$ is original dataset images, $c_{T}$ is user instruction, $s_{I}$ and $s_{T}$ are Classifier-Free-Guidance [20] scale for $c_{I}$ and $c_{T}$, $\epsilon_{\theta}(\tilde{I},c_{I},c_{T})$ is Coherent-IP2P denoiser networks including VAE [26]. This score-based guidance encourages a set of rendered 2D images from the pseudo-edited 4D Gaussians at arbitrary timesteps $\tilde{I}$ to resemble the edited images that IP2P would generate based on the $c_{I}$ and $c_{T}$, thereby effectively refining motion artifacts. As a result, we obtain refined canonical 3D Gaussians $\mathcal{G}_{\text{canon}}^{\text{ref}}$ that aligns well with the original deformation field $\left\{\mathcal{E}^{\text{opt}}(\mathcal{G}_{\text{canon}},t),\mathcal{D}^{\text{opt}}\right\}$ while maintaining the edited appearance. After the refinement stage, we obtain a completely edited 4D dynamic scene which is represented as $\left\{\mathcal{G}_{\text{canon}}^{\text{ref}},\mathcal{E}^{\text{opt}}(\mathcal{G}_{\text{canon}},t),\mathcal{D}^{\text{opt}}\right\}$.
@@ 5.1 Experimental Setup
We use DyNeRF [28] and Technicolor [50], a real-world multiview video dataset, to train and edit 4D dynamic scenes. The DyNeRF dataset includes six 10-second video sequences captured at 30 fps by 15 to 20 cameras with a face-forward perspective. Technicolor includes a wider variety of motion and scenarios, captured with 16 cameras. For comparison with the baseline [4], we trim the videos into 50-frame-long segments. We have also included the results on monocular datasets [41, 16] in the supplementary.
@@
We conduct a qualitative and quantitative comparison with Instruct 4D-to-4D [38], the only prior work addressing instruction-guided 4D dynamic scene editing. Instruct 4D-to-4D utilizes NeRFPlayer [55] as its backbone 4D representation and employs an iterative dataset update method, which involves editing all 2D images used for synthesizing the dynamic scene. It utilizes optical flow-based warping [58] and depth-based warping to ensure consistency across all edited 2D images. To alleviate the time-consuming dataset update process, Instruct 4D-to-4D employs two GPUs in parallel: one for the dataset update thread and the other for the dynamic scene editing thread.
@@
In Sec. 4.1, we follow the experimental settings of [63]. Throughout the experiments utilizing InstructPix2Pix [4], we set the CFG [20] scales for image condition and text instruction to 1.2 and 8.5 to 10.5, respectively. In the 3D Gaussian editing stage (Sec. 4.2), we train for 800 to 1000 iterations, depending on the editing style. For the score-based refinement stage (Sec. 4.3), an average of 800 iterations is sufficient to complete the dynamic scene editing successfully. All experiments are conducted using a single NVIDIA A40 GPU.
Figure 4: Effectiveness of score-based temporal refinement: Score-based temporal refinement effectively resolves misalignment between the canonical 3D Gaussians and the original deformation field that arises during the 3D Gaussian editing process. Without requiring additional 2D image updates, this process completes dynamic scene editing within a few hundred iterations.
@@ 5.2 Results
To quantitatively evaluate the visual quality of the edited dynamic scene, we measure PSNR, SSIM [61], and LPIPS [69] between the 2D multiview images used as supervision for dynamic scene editing and the images rendered from the edited dynamic scene using the corresponding camera parameters. Additionally, to assess how well the edited dynamic scene aligns with the input instruction, we also measure CLIP [45] similarity.
Table 1 presents a quantitative comparison of our method, Instruct-4DGS, and the baseline, Instruct 4D-to-4D, on DyNeRF. While our method shows slightly worse PSNR and SSIM in some cases, this is expected due to our efficient editing strategy. Unlike the baseline, which directly optimizes pixel-level accuracy using all edited images as training targets, our approach optimizes the dynamic scene using only instructions, without additional image editing during the temporal refinement stage. Consequently, pixel-wise accuracy (i.e., PSNR and SSIM) may be worse, but our method demonstrates superior perceptual quality, as shown in the consistently lower LPIPS across all cases. Additionally, our approach excels in instruction-following fidelity, achieving higher CLIP similarity than the baseline. Notably, our method accomplishes this 2–3 times faster while requiring fewer GPUs, making it significantly more efficient for real-world applications.
Table 2 compares the efficiency of the baseline and our method. In terms of editing time, our method completes editing 2–3 times faster while using only a single GPU, whereas the baseline requires two identical GPUs. This speed advantage could potentially become more pronounced as the number of timesteps in the dynamic scene increases. These results demonstrate that our method achieves efficiency by leveraging the static-dynamic separability of 4DGS and employing score-based temporal refinement, enabling significantly faster dynamic scene editing without extensive, time-consuming dataset updates.
@@
Ours 19.25 0.783 0.303 0.249
Table 1: Quantitative comparison of editing quality: Comparison of performance metrics between Instruct 4D-to-4D (I4D24D) and our Instruct-4DGS (Ours) under various editing instructions on DyNeRF. Higher values indicate better performance for PSNR, SSIM, and CLIP similarity; lower values are better for $\text{LPIPS}_{\text{VGG}}$.
@@
Instruct-4DGS 1 GPU 40 minutes
Table 2: Quantitative comparison of editing efficiency: Our proposed Instruct-4DGS significantly reduces editing time even with fewer GPU resources compared to the baseline.
@@
Figure 5: Qualitative results across various editing styles: Editing results of the scenes cook_spinach, flame_steak, and coffee_martini scenes from DyNeRF. Instruct-4DGS successfully edits dynamic scenes closely following the given user instructions.
@@
We conduct an ablation study to evaluate the impact of each design choice in our method, particularly their contributions to efficiency and quality. The qualitative and user study results are shown in Fig. 8. We recruited 50 participants of varying demographics, collecting a total of 50 preference rankings on the editing quality of videos generated by the four method variants.
@@
First, we examine dynamic scene editing using only score-based editing, without 3D Gaussian editing (denoted as “Fully SDS”). As shown in Fig. 8 (a), this approach preserves smooth motion but fails to ensure sufficient instruction alignment, leading to low-fidelity results. In contrast, incorporating the 3D Gaussian editing stage significantly improves fidelity while enabling effective motion refinement in the temporal refinement process. This highlights the importance of direct supervision via edited 2D images in maintaining fidelity and quality.
@@
To mitigate inherent issues of SDS such as Janus problem [53], and to provide stable guidance for the score-based temporal refinement, we employ Coherent-IP2P. To validate this choice, we compare results by refining the pseudo-edited dynamic scene with the original IP2P (denoted as “Refine w/ original IP2P”). As shown in Fig. 8 (b), using the original IP2P for temporal refinement leads to severe visual artifacts and low-quality outputs, whereas Coherent-IP2P preserves details and retains the scene’s semantics. This confirms that Coherent-IP2P mitigates noisy guidance and blurry artifacts by enabling information sharing among images within the same batch.
Figure 6: Qualitative comparison of visual quality: We compare our method with the baseline [38] on DyNeRF [28] coffee_martini and sear_steak scenes, as well as Technicolor [50]’s Painter and Train scenes. See supplementary for more results.
@@
Figure 7: Qualitative comparison of temporal consistency: The baseline shows noticeable flickering artifacts across timesteps. In contrast, Instruct-4DGS effectively avoids such artifacts by editing only the static component with score-based temporal refinement.
@@
Figure 8: Ablation study of the dynamic scene editing method: Each pie chart shows the proportion of user preferences (1st-4th ranks) for each method variant. Our proposed method (denoted as “Ours (w/o refine $\{\mathcal{E},\mathcal{D}\}$)”) achieves the highest preference score.
Lastly, to evaluate the effectiveness of refining the deformation field, we compare our final method (denoted as “Ours (w/o refine $\{\mathcal{E},\mathcal{D}\}$)”)—which refines only the static 3D Gaussians, excluding the deformation field—with “Ours (w/ refine $\{\mathcal{E},\mathcal{D}\}$)”. As shown in Fig. 8 (c), refining the deformation field introduces temporal inconsistencies and motion artifacts. In contrast, our final method effectively preserves temporal coherence while maintaining high editing fidelity. These results indicate that refining the deformation field does not contribute positively to dynamic scene editing and can instead introduce undesirable distortions.
6 Conclusion and Limitations